OpenStudy (anonymous):

2x + y = 7 5x + y = 9 Solve the system of equations.

OpenStudy (anonymous):

yo its easy where are you having problems? let;s solve first equation for y what do you get?

OpenStudy (anonymous):

ヽ༼ຈل͜ຈ༽ノ

OpenStudy (hihi67):

Okay 2x+y=7 Divide by 2 x+y=3.5

OpenStudy (anonymous):

im still confused

OpenStudy (hihi67):

Number 2 5x+y=9 Divide by 5 x+y=1.8

OpenStudy (hihi67):

What grade are you in?

OpenStudy (anonymous):

12th

OpenStudy (anonymous):

i hate math so much

OpenStudy (hihi67):

Okay, so senior in high school?

OpenStudy (anonymous):

its frustrating

OpenStudy (anonymous):

yes i am

OpenStudy (hihi67):

Okay, cool.

OpenStudy (hihi67):

Well, do you know how to solve an equation?

OpenStudy (anonymous):

yes

OpenStudy (hihi67):

Well, this is nothing more than an equation.

OpenStudy (anonymous):

its easy from first equation we get y=7-2x pby putting this value in second equation we get 5x+7-2x=9 3x+7=9 x=2/3

OpenStudy (anonymous):

you can solve this also by elimination which means you add the two equations 2x + y = 7 +5x + y = 9 <---make this equation negative in order to subtract to eliminate the y 2x + y = 7 -5x - y = -9 -3x=-2 divide by -3 x=2/3 then input x into either equation i picked the first 2(2/3)+y=7 4/3+y=7(which is 21/3) subtract 4/3 y=17/3